OpenStudy (anonymous):

Find the equation of the line that passes through the pair of points. Write your answer in slope-intercept form. (5, -6\), (-5, 0\)

OpenStudy (anonymous):

do i do the y2-y1/x2-x1 ?

OpenStudy (saifoo.khan):

yes, that's the first step.

OpenStudy (anonymous):

That will give you the slope

OpenStudy (anonymous):

ok brb 30 seconds. im gonna do that first

OpenStudy (anonymous):

-6?

OpenStudy (saifoo.khan):

-0.6

OpenStudy (anonymous):

dude your making me lag really bad when you do that rofl

OpenStudy (anonymous):

y=mx+b the "m" is what you are solving for with the slope, the "b" is the y-intercept

OpenStudy (anonymous):

saifoo, go help some junior high kids

OpenStudy (saifoo.khan):

why the heck is that? o_O

OpenStudy (anonymous):

i had 0-6/-5-5. ws that wrong?

OpenStudy (anonymous):

Because your spamming

OpenStudy (anonymous):

Just chill

OpenStudy (anonymous):

1 sec on the slope

OpenStudy (saifoo.khan):

\[\Huge \frac{0+6}{-5-5}= \frac{6}{-10}=-0.6\]

OpenStudy (anonymous):

ya saifoo is right

OpenStudy (anonymous):

I actually did this stuff like 6 years ago so I'm not the most fresh on it, sean knows what hes doing, listen to him.

OpenStudy (anonymous):

but the equation has -6. take another look at it

OpenStudy (anonymous):

(0-(-6))/(-5-5) is the slope intercept formula. This is your m in the equation y=mx+b. The b is your y-intercept. After you solve for m, you can plug in either set of (x,y) values and get your b. Hope this helps! Thanks, Sean Walsh Facebook: Tutor Sean http://www.tutorsean.net

OpenStudy (saifoo.khan):

where is that equation u r talking about? @dude.

OpenStudy (anonymous):

its 0 - (-6) subtracting a negative gives a positive value

OpenStudy (anonymous):

ah. ok. derp

OpenStudy (anonymous):

Sean do you know laplace transforms at all?

OpenStudy (anonymous):

So your answer would be y = -6/10x + b. If you want b, then you would have: 0 = -6/10(-5)+b Which evaluates to: 0 = 30/10 + b 0 = 3 + b -3 = b So final equation is: y = -6/10x - 3

OpenStudy (anonymous):

@Pan, no. Unfortunately, I am a computer science major and Diff Eq is not required, but I am taking next term. :(

OpenStudy (anonymous):

Ah ok, no worries

OpenStudy (anonymous):

At least you know what it is, better than I could say before I took DEQ

OpenStudy (anonymous):

I only know because I have many friends who are math profs. Hear that they are EXTREMELY useful though. Has something to do with integrals right? if you like, go ahead and like Tutor Sean on Facebook. I could use someone on there who knows Diff Eq if you are interested. It is another community where I help people solve math problems. I could make you an admin so that you can answer questions. :)